# Impact of reactor antineutrinos on the neutrino floor in low-mass WIMP-like dark matter searches

### v1

@@ Abstract
The sensitivity of conventional direct dark matter searches for weakly interacting massive particles (WIMPs) is ultimately limited by coherent elastic neutrino–nucleus scattering (CENS), which produces nuclear recoils indistinguishable from WIMP signals and defines the so-called neutrino floor. While the effects of solar, geoneutrino, diffuse supernova, and atmospheric neutrinos have been extensively studied in this context, the contribution from reactor antineutrinos has received comparatively little attention. We present the first systematic evaluation of how reactor antineutrino fluxes, modeled as a function of reactor–detector distance, modify the neutrino floor for low-mass WIMP searches using SuperCDMS-like high-voltage germanium detectors. Both discovery-limit and opacity-based formulations of the neutrino floor are examined under consistent assumptions. We find that proximity to gigawatt-scale reactors within 10 km can raise the neutrino floor by up to a few orders of magnitude, significantly reducing the sensitivity to sub-10 GeV/ dark matter. Beyond 100 km, the reactor contribution becomes negligible. These results establish reactor proximity as a critical site-dependent parameter in defining the neutrino floor and provide guidance for optimizing the design and siting of future low-threshold dark matter experiments.
@@ I Introduction
The existence of dark matter (DM) is supported by a wide range of astrophysical and cosmological observations first_proposed_DM:1933gu; rotation_curve_1970; bullet_cluster_2006; Planck_2015_CMB. However, its particle nature remains unknown. Weakly Interacting Massive Particles (WIMPs) are theoretically well-motivated and represent one of the leading dark matter candidates WIMP_1_1984; WIMP_2_1995_supersymmetric; Relic_WIMP_abundances_2012. Direct detection experiments search for nuclear recoils produced by the elastic scattering of WIMPs off target nuclei. Since both the WIMP mass () and the WIMP–nucleon scattering cross section () are unknown, these experiments explore a wide parameter space in and CDMSliteR3_WIMP_PLR_2019; CRESST_2024_WIMP; LZ_2023_WIMP; PandaX_2025_WIMP; XENONnT_2025_WIMP.
@@ II.3 Quenching and the Measured WIMP Energy Spectrum
where and is the dimensionless energy parameter. The constant is material-dependent. For germanium (), recent CENS results from the CONUS+ Collaboration CONUS_reactor_anti-neutrino_observation_Ackermann:2025obx indicate that provides good agreement down to recoil energies of . We adopt this value and assume that the Lindhard model remains valid at lower recoil energies. Because both WIMP- and CENS-induced interactions produce nuclear recoils, they are expected to exhibit similar quenching behavior. Consequently, the exact choice of quenching model has little impact on the determination of the neutrino floor. However, applying quenching is important for obtaining a realistic prediction of the experimentally observable recoil spectrum.
@@ III.1 Differential Scattering Cross Section
Coherent elastic neutrino–nucleus scattering (CENS) occurs when a neutrino scatters elastically off an entire nucleus at low momentum transfer, resulting in a coherent enhancement of the cross section proportional to the square of the weak nuclear charge freedman. Within the Standard Model, the differential cross section for a neutrino of energy scattering off a nucleus of mass and producing a recoil energy can be expressed as CEvNS_Crosssection_Scholberg:2005qs:
@@ III.3.1 Non-Reactor Neutrino Sources
Geoneutrinos originate from the -decays of long-lived isotopes such as uranium, thorium, and potassium within the Earth’s crust and mantle. Their flux exhibits moderate spatial variation due to differences in crustal composition and thickness. The geoneutrino flux is typically two orders of magnitude lower than the solar neutrino flux; therefore, its contribution to the CENS background is subdominant. Geoneutrinos produce low-energy recoils, primarily affecting WIMP sensitivity near GeV/. Given the moderate site-to-site variations and the overall lower flux compared to solar neutrinos, we adopt the differential CENS event rate in germanium detectors in terms of nuclear recoil energy evaluated for the China Jinping Underground Laboratory (CJPL) as a representative case Geo_neutrino_Gelmini:2018gqa.
@@
Solar neutrinos dominate the CENS background at low WIMP masses (), while atmospheric neutrinos dominate at higher masses (). The fractional uncertainties associated with each neutrino source are summarized in Table 1. The uncertainty in the CONUS-reported quenching parameter is on the order of 1% COUNSplus_CEvNS, which is much smaller than the overall systematic uncertainties in the CENS rate. Since both CENS and WIMP-induced spectra are subject to similar quenching effects, the impact of quenching uncertainty on the neutrino floor determination is expected to be minimal and is therefore neglected in our calculations.
@@
| pp | 0.6% |
| pep | 1% |
| hep | 30% |
| 7Be I (384 keV) | 6% |
| 7Be II (861 keV) | 6% |
| 8B | 2% |
| 13N | 15% |
| 15O | 17% |
| 17F | 20% |
| DSNB | 50% |
| Atmospheric | 25% |
| Geo (40K) | 6% |
| Geo (238U) | 30% |
| Geo (232Th) | 30% |
| Reactor | 8% |
@@ III.3.2 Reactor Antineutrinos
At low energies (around 1 MeV and below), an additional contribution to the flux arises from neutron capture on 238U, with the corresponding spectrum taken from Ref. TEXONO_U238_capture:2006xds. This spectrum was normalized and scaled according to its antineutrino yield per fission, taken as 1.2 CONNIE_anti_neutrino_spectrum:2019xid. The combined reactor antineutrino spectra at different reactor–detector distances are shown in Fig. 2.
Reactor antineutrinos can constitute an important component of the CENS background for direct detection WIMP search experiments, especially for sites located near nuclear power plants. Their contribution strongly depends on the distance between the detector and the reactor core, leading to large variations in the expected event rates. As shown in Fig. 3, the CENS event rate can vary by a few orders of magnitude for detector locations within 50 km of a reactor, surpassing the solar neutrino background at distances below 10 km. Accurate modeling of the site-dependent reactor flux is therefore essential for evaluating the neutrino floor and assessing the discovery reach of future low-threshold dark matter experiments.
We validated our reactor antineutrino CENS recoil spectrum against the CENS spectrum measured by the CONUS+ experiment CONUS+_first_observation_2025. As discussed in Appendix A, our Standard Model prediction for the CENS recoil spectrum is in good agreement with the experimental data from CONUS+, measured at the same KKL reactor.
@@ IV.1 Discovery-Limit Based Neutrino Floor
In the discovery-limit framework Billard:2013qya, the neutrino floor is defined as the minimum WIMP–nucleon cross section for which a WIMP signal can be detected with at least significance in 90% of hypothetical experiments, in the presence of CENS backgrounds.
@@
The exposures used in the pseudo-experiments are chosen to provide statistically meaningful sampling of the neutrino background. For the low-threshold setup, the exposure corresponds to roughly 200 expected 8B solar neutrino events, while the high-threshold configuration yields about 500 total neutrino events Ruppin_followup:2014bra; Billard:2013qya. These correspond to exposures of approximately 0.4 ton-year and ton-year, respectively. With these choices, the statistical uncertainty on the total neutrino background remains smaller than the systematic uncertainty in the neutrino flux, allowing the pseudo-experiments to capture both statistical fluctuations and systematic uncertainties. It should be noted that these threshold and exposure values are somewhat idealized and not yet representative of current detector capabilities. Prototype detectors such as those developed by SuperCDMS HVeVR1; SuperCDMS:HVeV_compton_steps; SuperCDMS:HVeV4_DM and CRESST CRESST_2024_WIMP have achieved eV-scale thresholds but remain limited to gram-scale target masses, making ton-year exposures currently unfeasible.
@@ IV.2 Opacity-Based Neutrino Floor
Appendix B presents a comparison of the baseline neutrino floor—i.e., the neutrino floor obtained without including reactor antineutrinos—from this study with previously published results. The comparison is shown for both the discovery-limit–based approach Billard:APPEC_report and the opacity-based approach OHare:2021utq, focusing on Ge-based detectors. Overall, our baseline limits show reasonable agreement with previously published results under both definitions.
@@ V Implications for Future Dark Matter Experiments
Although this effect is negligible for deep underground laboratories such as SNOLAB—where typical reactor baselines exceed 100 km reactor_anti_neutrino_contribution_PRD_2024—it becomes a vital consideration for next-generation low-threshold experiments and site planning. Accurate inclusion of reactor antineutrino fluxes in sensitivity projections is therefore essential for realistic assessments of discovery reach in the sub-10 GeV/ WIMP mass regime. Mitigation strategies, including directional detection Directional_search_OHare:2015utx; Directional_search_Grothaus:2014hja, could further help preserve sensitivity in reactor-proximate environments.
VI Summary and Conclusions
@@ Acknowledgments
Appendix A Validation of reactor CENS differential rate
@@
Appendix B Comparison of baseline neutrino floor

### v2

@@ Abstract
The sensitivity of conventional direct dark matter searches for weakly interacting massive particles (WIMPs) is ultimately limited by coherent elastic neutrino–nucleus scattering (CENS), which produces nuclear recoils indistinguishable from WIMP signals and defines the so-called neutrino floor. While the effects of solar neutrinos, geoneutrinos, diffuse supernova neutrinos, and atmospheric neutrinos have been extensively studied in this context, the contribution from reactor antineutrinos has received comparatively little attention. We present the first systematic evaluation of how reactor antineutrino fluxes, modeled as a function of reactor–detector distance, modify the neutrino floor for low-mass WIMP searches using SuperCDMS-like high-voltage germanium detectors. Both discovery-limit and opacity-based formulations of the neutrino floor are examined under consistent assumptions. We find that proximity to gigawatt-scale reactors within 10 km can raise the neutrino floor by up to a few orders of magnitude, significantly reducing the sensitivity to sub-10 GeV/ dark matter. Beyond 100 km, the reactor contribution becomes negligible. These conclusions hold for both definitions of the neutrino floor and remain stable under reasonable variations in detector quenching, site-dependent geoneutrino flux, and reactor antineutrino flux uncertainties, emphasizing reactor proximity as a critical factor in site selection for future low-threshold dark matter experiments.
@@ I Introduction
The existence of dark matter (DM) is supported by a wide range of astrophysical and cosmological observations first_proposed_DM:1933gu; rotation_curve_1970; bullet_cluster_2006; Planck_2015_CMB. However, its particle nature remains unknown. Weakly Interacting Massive Particles (WIMPs) are theoretically well-motivated and represent one of the leading dark matter candidates WIMP_1_1984; WIMP_2_1995_supersymmetric; Relic_WIMP_abundances_2012. Direct detection experiments search for nuclear recoils produced by the elastic scattering of WIMPs off target nuclei. Since both the WIMP mass () and the WIMP-nucleon scattering cross section () are unknown, these experiments explore a wide parameter space in and CDMSliteR3_WIMP_PLR_2019; CRESST_2024_WIMP; LZ_2023_WIMP; PandaX_2025_WIMP; XENONnT_2025_WIMP.
@@ II.3 Quenching and the Measured WIMP Energy Spectrum
where and is the dimensionless energy parameter. The constant is material-dependent. For germanium (), recent CENS results from the CONUS+ Collaboration CONUS_reactor_anti-neutrino_observation_Ackermann:2025obx indicate that provides good agreement down to recoil energies of . We adopt this value and extrapolate it below , assuming that the Lindhard model remains valid in this regime. Because both WIMP- and CENS-induced interactions produce nuclear recoils, they are expected to exhibit similar quenching behavior. Consequently, the exact choice of quenching model has little impact on the determination of the neutrino floor. However, applying quenching is important for obtaining a realistic prediction of the experimentally observable recoil spectrum.
@@ III.1 Differential Scattering Cross Section
Coherent elastic neutrino–nucleus scattering (CENS) occurs when a neutrino scatters elastically off an entire nucleus at low momentum transfer, resulting in a coherent enhancement of the cross section proportional to the square of the weak nuclear charge freedman. The process is neutrino–flavor independent, as it is mediated by the neutral-current interaction. Within the Standard Model, the differential cross section for a neutrino of energy scattering off a nucleus of mass and producing a recoil energy can be expressed as CEvNS_Crosssection_Scholberg:2005qs:
@@ III.3.1 Non-Reactor Neutrino Sources
Geoneutrinos originate from the -decays of long-lived isotopes such as uranium, thorium, and potassium within the Earth’s crust and mantle. Their flux exhibits moderate spatial variation due to differences in crustal composition and thickness. The geoneutrino flux is typically two orders of magnitude lower than the solar neutrino flux; therefore, its contribution to the CENS background is subdominant. Geoneutrinos produce low-energy recoils, primarily affecting WIMP sensitivity near GeV/. Given the moderate site-to-site variations and the overall lower flux compared to solar neutrinos, we adopt the differential CENS event rate in germanium detectors in terms of nuclear recoil energy evaluated for the China Jinping Underground Laboratory (CJPL) as a representative case Geo_neutrino_Gelmini:2018gqa. Appendix A shows that reducing the geoneutrino flux by a factor of two, roughly representing SNOLAB levels geo_neutrino_SNOLAB_vs_CJPL; Geo_neutrino_Gelmini:2018gqa, modifies the no-reactor neutrino floor contributions by at most 10% near GeV/. Therefore, the reactor antineutrino flux, which scales with the reactor–detector distance, quickly dominates site-dependent variations, and a single representative geoneutrino model suffices for this study.
@@
Solar neutrinos dominate the CENS background at low WIMP masses (), while atmospheric neutrinos dominate at higher masses (). The fractional uncertainties associated with each neutrino source are summarized in Table 1. The uncertainty in the CONUS+ reported quenching parameter (=0.162 0.004) is on the order of 1% COUNSplus_CEvNS, which is much smaller than the overall systematic uncertainties in the CENS rate. Since both CENS and WIMP-induced spectra are subject to similar quenching effects, the impact of quenching uncertainty on the neutrino floor determination is expected to be minimal. We evaluated this systematic impact by varying within its uncertainty range and found that the resulting neutrino floor remain effectively unchanged. Consequently, we neglect the propagation of quenching-parameter uncertainty in our main calculations. Further discussion is provided in Appendix B.
@@
| pp | 0.6% Overcome_neutrino_floor:2020lva |
| pep | 1% Overcome_neutrino_floor:2020lva |
| hep | 30% Overcome_neutrino_floor:2020lva |
| 7Be I (384 keV) | 6% Overcome_neutrino_floor:2020lva |
| 7Be II (861 keV) | 6% Overcome_neutrino_floor:2020lva |
| 8B | 2% Overcome_neutrino_floor:2020lva |
| 13N | 15% Overcome_neutrino_floor:2020lva |
| 15O | 17% Overcome_neutrino_floor:2020lva |
| 17F | 20% Overcome_neutrino_floor:2020lva |
| DSNB | 50% Overcome_neutrino_floor:2020lva |
| Atmospheric | 25% Overcome_neutrino_floor:2020lva |
| Geo (40K) | 6% Geo_neutrino_Gelmini:2018gqa |
| Geo (238U) | 30% Geo_neutrino_Gelmini:2018gqa |
| Geo (232Th) | 30% Geo_neutrino_Gelmini:2018gqa |
| Reactor | 8% Overcome_neutrino_floor:2020lva |
@@ III.3.2 Reactor Antineutrinos
At low energies (around 1 MeV and below), an additional contribution to the flux arises from neutron capture on 238U, with the corresponding spectrum taken from Ref. TEXONO_U238_capture:2006xds. This spectrum was normalized and scaled according to its antineutrino yield per fission, taken as 1.2 CONNIE_anti_neutrino_spectrum:2019xid. The combined reactor antineutrino spectra at different reactor–detector distances are shown in Fig. 2. Following Ref. Overcome_neutrino_floor:2020lva, we adopt a nominal reactor antineutrino flux uncertainty of 8%. This value is broadly in line with the 7.8% deficit reported by Daya Bay in the observed versus predicted yield of the 235U fission isotope DayaBay_reactor_flux_uncertainty.
Reactor antineutrinos can constitute an important component of the CENS background for direct detection WIMP search experiments, especially for sites located near nuclear power plants. Their contribution strongly depends on the distance between the detector and the reactor core, leading to large variations in the expected event rates. As shown in Fig. 3, the CENS event rate can vary by a few orders of magnitude for detector locations within 50 km of a reactor and begins to surpass the solar neutrino background at baselines below 10 km. Accurate modeling of the site-dependent reactor flux is therefore essential for evaluating the neutrino floor and assessing the discovery reach of future low-threshold dark matter experiments.
We validated our reactor antineutrino CENS recoil spectrum against the CENS spectrum measured by the CONUS+ experiment CONUS+_first_observation_2025. As discussed in Appendix C, our Standard Model prediction for the CENS recoil spectrum is in good agreement with the experimental data from CONUS+, measured at the same KKL reactor.
@@ IV.1 Discovery-Limit Based Neutrino Floor
In the discovery-limit framework Billard:2013qya, the neutrino floor is defined as the minimum WIMP-nucleon cross section for which a WIMP signal can be detected with at least significance in 90% of hypothetical experiments, in the presence of CENS backgrounds.
@@
The exposures used in the pseudo-experiments are chosen to provide statistically meaningful sampling of the neutrino background. For the low-threshold setup, the exposure corresponds to roughly 200 expected 8B solar neutrino events, while the high-threshold configuration yields about 500 total neutrino events Ruppin_followup:2014bra; Billard:2013qya. These correspond to exposures of approximately 0.4 ton-year and ton-year, respectively. With these choices, the statistical uncertainty on the total neutrino background remains smaller than the systematic uncertainty in the neutrino flux, allowing the pseudo-experiments to capture both statistical fluctuations and systematic uncertainties.
It should be noted that these ultra-low threshold and ultra-high exposure values are somewhat idealized and not yet representative of current detector capabilities. They serve only as mathematical constructs that allow the neutrino floor to be evaluated in the asymptotic, systematics-dominated regime. Prototype detectors such as those developed by SuperCDMS HVeVR1; SuperCDMS:HVeV_compton_steps; SuperCDMS:HVeV4_DM and CRESST CRESST_2024_WIMP have achieved eV-scale thresholds but remain limited to gram-scale target masses, making ton-year exposures currently unfeasible.
@@ IV.2 Opacity-Based Neutrino Floor
Appendix D presents a comparison of the baseline neutrino floor, i.e., the neutrino floor obtained without including reactor antineutrinos, from this study with previously published results. The comparison is shown for both the discovery-limit–based approach Billard:APPEC_report and the opacity-based approach OHare:2021utq, focusing on Ge-based detectors. Overall, our baseline limits show reasonable agreement with previously published results under both definitions.
@@ V Implications for Future Dark Matter Experiments
We also note that future experiments may be exposed to antineutrino fluxes originating from multi-reactor clusters, where individual reactors can have different thermal powers and baselines. In such environments, the total CENS background is simply the sum of the contributions from each reactor, weighted by their respective powers and source–detector distances. As discussed in Appendix E, a representative configuration consisting of reactors with 1.2 GWth at 5 km, 2.4 GWth at 10 km, and 3.6 GWth at 50 km yields a combined antineutrino spectrum that is effectively equivalent to that from a single-reactor configuration with an effective baseline of 7 km and a total thermal power of 3.6 GWth. Consequently, the resulting neutrino floor is nearly indistinguishable from that of this equivalent single-reactor model. Similarly, annual duty-cycle variations in reactor operation can be interpreted as a reduction in the effective thermal power, leading to a proportionally reduced antineutrino flux. Finally, variations in the reactor antineutrino spectrum arising from changes in reactor fuel fission fractions have only a minor impact. As discussed in Appendix F, varying the fission composition modifies the predicted neutrino floor by at most 10% at a reactor to detector distance of 5 km, far smaller than the order-of-magnitude enhancement of the neutrino floor relative to the no-reactor scenario.
@@
Although the impact of reactor antineutrino would be negligible for deep underground laboratories such as SNOLAB, where typical reactor baselines exceed 100 km reactor_anti_neutrino_contribution_PRD_2024, it becomes a vital consideration for next-generation low-threshold experiments and site planning. Accurate inclusion of reactor antineutrino fluxes in sensitivity projections is therefore essential for realistic assessments of discovery reach in the sub-10 GeV/ WIMP mass regime. Mitigation strategies, including directional detection Directional_search_OHare:2015utx; Directional_search_Grothaus:2014hja, could further help preserve sensitivity in reactor-proximate environments.
VI Summary and Outlook
@@
Studies incorporating reasonable variations in the detector’s quenching response, geoneutrino contribution, and uncertainties in the reactor antineutrino flux show that the dominant factor driving the loss of discovery potential is the geometric scaling of the reactor antineutrino flux.
@@
Recent work has shown that neutrino-induced Migdal events can generate an additional “neutrino fog” at sub-GeV/c2 masses, where both CENS and WIMP interactions may trigger Migdal electron emission or excitation migdal_neutrino_floor. Such effects are especially relevant for inelastic dark-matter search channels SuperCDMS_migdal_Brem_WIMP_2023. In the present study, however, we restrict ourselves to the standard formulation of the neutrino floor based solely on elastic nuclear recoils, consistent with the original definitions in Refs. Billard:2013qya; OHare:2021utq and widely adopted in the literature. A dedicated investigation of how reactor-dependent backgrounds modify the Migdal-induced neutrino fog would require a distinct treatment of both the WIMP signal and CENS background models, and is beyond the scope of this work. This represents a natural direction for future study.
@@ Acknowledgments
Appendix A Impact of Geoneutrino Flux Variations on the Neutrino Floor
We adopt the geoneutrino flux from the China Jinping Underground Laboratory (CJPL) as a reference, which is relatively high due to the surrounding crustal composition. In comparison, SNOLAB, planned for next-generation SuperCDMS experiments, has a lower crustal geoneutrino flux, approximately half that of CJPL, reflecting its distinct local crustal and geological environment geo_neutrino_SNOLAB_vs_CJPL.
To quantify the impact of site-dependent geoneutrino flux uncertainties, we performed a variation study by scaling the geoneutrino contribution by a factor of two, in the absence of reactor neutrino contributions. As shown in Fig. 10, this variation modifies the neutrino floor by up to 10% near . Overall, the geoneutrino contribution remains subdominant compared to the solar neutrino background.
For detectors exposed to a gigawatt-scale reactor at 10 km, the geoneutrino effect becomes subdominant, quickly suppressed by the reactor antineutrino flux. This demonstrates that while local crustal composition can slightly affect the neutrino floor, reactor proximity dominates for near-reactor sites.
Appendix B Effect of Quenching on the Neutrino Floor
To evaluate the effect of quenching-model uncertainties on the neutrino floor, we studied the no-reactor scenario as a representative benchmark. The analysis uses the Lindhard parametrization of the ionization yield, with the quenching parameter reported by CONUS+, COUNSplus_CEvNS. To assess the sensitivity of the neutrino floor to this uncertainty, we varied the quenching parameter within its range. This variation modifies both the magnitude and the shape of the corresponding Lindhard quenching curve.
Figure 11 shows the resulting neutrino floors obtained for the nominal, , and cases. The resulting neutrino floors are consistent across the full WIMP-mass range, demonstrating that the neutrino floor is robust against reasonable variations in the quenching parameter. This behavior is consistent with the expectation that both CENS and WIMP-induced spectra are modified in a similar manner by this detector effect.
Although the quenching uncertainty has a negligible impact on the neutrino-floor calculation itself, a realistic quenching model remains essential for predicting measurable recoil spectra and for performing spectral analyses with experimental data.
Appendix C Validation of reactor CENS differential rate
@@
Appendix D Comparison of baseline neutrino floor
@@
Appendix E Influence of Multiple Nearby Reactors on the Neutrino Floor
Experiments often encounter antineutrino fluxes originating from multiple reactors with different thermal powers and baselines. To evaluate the implications of such multi-reactor environments for our study, we computed the neutrino floor for a representative mixed-reactor configuration in which the total antineutrino flux at the detector is obtained by summing the contributions from all reactors. Each contribution is weighted by its thermal power and by the inverse square of the source-to-detector distance, producing a combined CENS spectrum that reflects the full cluster.
The specific configuration considered consists of a 1.2 GW reactor at 5 km, a 2.4 GW reactor at 10 km, and a 3.6 GW reactor at 50 km. Figure 14 shows the neutrino floor obtained for this mixed configuration. The resulting total antineutrino spectrum is equivalent to that produced by a single 3.6 GW reactor located at an effective distance of approximately 7 km. Consequently, the neutrino floor obtained for the mixed-reactor case is effectively identical to that of the corresponding single-reactor effective configuration.
This demonstrates that, for the purposes of neutrino-floor estimation, the mixed-reactor configuration considered here is effectively equivalent to a single-reactor scenario with comparable flux normalization. Similarly, annual duty-cycle averaging of reactor operations can lead to a reduced effective antineutrino flux, which may be treated as a single-reactor configuration with a correspondingly lower thermal power.
Appendix F Impact of Reactor Antineutrino Spectrum Shape Variations
Nearby reactors may operate with different fuel compositions, leading to variations in their fission fractions and, consequently, in the shape of the emitted antineutrino spectrum. As discussed in Sec. III.3.2, changes in the relative contributions of 235U, 238U, 239Pu, and 241Pu affect both the normalization and energy dependence of the flux. To quantify this effect, we consider two representative configurations. The first corresponds to a commercial light-water reactor with fission fractions conus+_experiment_2024. The second represents a highly enriched research reactor such as HFIR, in which 235U accounts for approximately 99% of all fissions HFIR_fission_frac.
For both scenarios, we compute the CENS event rate assuming a thermal power of 3.6 GW and a reactor–detector separation of 5 km. The resulting variation in the neutrino floor is at the level of 10%, which is well below the order-of-magnitude enhancement observed at 5 km relative to the no-reactor baseline. Therefore, uncertainties in reactor fuel composition do not constitute a dominant systematic uncertainty in the determination of the neutrino floor.